A uniform plank of length 2.00 m and mass 26.0 kg
Answer #1

Equating the Vertical Forces,
T1*sin(40) + T2 = 26.0 * 9.8 + 715 -------1

Equating the Horizontal Forces,
T1 * cos(40) = T3 ------2

Taking the moment about the Left end,
715 * d + 26.0*9.8* (2.0/2) = T1 * sin(40) * 2.0
715 * 0.5 + 26.0*9.8* (2.0/2) = T1 * sin(40) * 2.0
T1 = 476.3 N

476.3*sin(40) + T2 = 26.0 * 9.8 + 715
T2 = 663.6 N

T3 = 476.3 * cos(40)
T3 = 364.9 N